OpenStudy (anonymous):

I already answered most of this question, I just need help with the rest. Write the expression x2 + 2x - 24 as a product of two linear expressions. Show your work and justify each step.

OpenStudy (anonymous):

x^2 + 2x - 24 GCF = 1 1 * -24 = -24 Factor of -24 that adds to give 2 = 1 x^2 + (1x + 1x) - 24 x^2 + 1x +1x - 24 x(x + 1) (x - 24)

OpenStudy (anonymous):

I answered most of the question, but I just need help with the rest please

OpenStudy (mathmate):

``` Factor of -24 that adds to give 2 = 1 ``` You need the two factors which multiply to -24 and add to 2. 1+1=2, but 1*1\(\ne\)24 You may want to rework from here.

OpenStudy (anonymous):

x^2 + 2x - 24 =x^2+6x-4x-24 =x^2-4x+6x-24 =x(x-4)+6(x-4) =(x+6)(x-4)